# Team RAS in 9th ABAW Competition: Multimodal Compound Expression Recognition Approach

### v1

@@ 1 Introduction
Our approach introduces a novel multimodal framework specifically designed for CER. It integrates six diverse modalities – static and dynamic facial expressions, audio signals, scene context, textual descriptions, scene-label alignment, probability and feature fusion – into a unified system. In contrast to prior work that relies primarily on static facial features [32, 34, 18], our method incorporates temporal dynamics using Transformer- and Mamba-based architectures, enabling more accurate modeling of affective evolution over time.
Moreover, unlike systems that require extensive fine-tuning on labeled CE corpora [22, 20, 24], our pipeline leverages zero-shot-capable components. These include Contrastive Language-Image Pretraining (CLIP)-driven label matching and semantic scene analysis using Qwen-VL, allowing the model to generalize beyond predefined emotion classes without domain-specific retraining. We also propose a Multi-Head Probability Fusion (MHPF) model that adaptively balances contributions from individual modality-specific predictors.
@@ 2 Related Work
We compare our proposed method with the State-of-the-Art (SOTA) methods presented in the scope of the 6th, 7th, 8th Affective Behavior Analysis in-the-Wild (ABAW) Competitions [14]. This challenge has seen the development of several CER systems, including our own [27].
Savchenko [28, 29] developed a series of methods for emotion recognition using both audio-visual and purely visual inputs. In the first method [28], the Wav2Vec2 model [1] was used to extract features from raw audio signals, facial features were extracted using both convolutional and transformer-based models trained on emotion recognition corpora. CER was carried out through score-level fusion and feature clustering. In a follow-up study [29], a visual-only method was proposed using lightweight EmotiNet models pre-trained for basic facial expression recognition. CE predictions were obtained by aggregating the outputs of basic emotions and applying temporal smoothing with box and Gaussian filters. This method enhanced stability while maintaining computational efficiency, making it applicable to real-time scenarios.
Yu et al. [34] and Wang et al. [32] proposed visual methods based on static models trained specifically for CER. In both methods, the Real-world Affective Faces Database (RAF-DB) annotated with CEs was used for model training. Notably, Wang et al. [32] also employed a visual-language method using the pre-trained Claude3222https://www.anthropic.com/claude model to generate additional annotations for test samples in the C-EXPR-DB [17].
Qiu et al. [22] developed an ensemble audio-visual method based on a dynamic CER model, distinguishing it from previous static methods. Convolutional features were extracted from audio spectrograms and facial regions, concatenated, and passed through transformer layers. Final CE predictions were obtained via majority voting across multiple models. The training process relied on a proprietary corpus annotated with CEs.
Lu et al. [20] proposed a vision-language method for CER, adapting the large-scale Qwen-VL model through a two-stage LoRA fine-tuning pipeline. In the first stage, the model is trained to recognize basic emotional representations, while the second stage focuses on refining the model using CE features. The method relies on prompt engineering to direct the model’s attention toward semantically relevant facial regions and emotion-related concepts.
@@
In the methods proposed by [22, 20, 18], additional corpora annotated with CEs were used to train models. Therefore, our method is more comparable to those of [32, 34, 27, 26, 24, 28, 29, 19], which rely on pretrained components and avoid direct training on the target corpus, thus operating in a zero-shot or lightly adapted setting.
@@ 3.1 Face Models
For each segment, the frame rate is downsampled to 20 frames per 4-second window, ensuring uniform temporal resolution and reducing computational complexity. Facial detection is performed using a two-stage method: an initial YOLO v11333https://github.com/akanametov/yolo-face model, pre-trained on WIDER FACE, and fine-tuned for real-world conditions, detects faces under challenging conditions, while a secondary refinement step based on MediaPipe FaceMesh [21] validates face geometry and filters out false positives detections.
@@ 3.6 Modality Fusion Models
In functional mapping, $\mathcal{P}=\{P^{(1)},\dots,P^{(M)}\}$ is a set of input probability distributions from $M$ models, where each $P^{(m)}\in\Delta^{C}$ lies on the $C$-dimensional simplex. For each attention head $h\in\{1,\dots,H\}$, class-specific weights are computed as:
$W_{h}^{(m)}=\text{Softmax}_{m}(\mathbf{H}_{w}^{(h,c)})\quad\forall c\in\{1,\dots,C\}$ (1)
$O_{h}=\sum_{m=1}^{M}W_{h}^{(m)}\odot P^{(m)}$ (2)
where $\mathbf{H}_{w}\in\mathbb{R}^{H\times M\times C}$ is the learnable weight tensor, and $\odot$ denotes element-wise multiplication. The outputs $O_{h}$ are then combined using softmax-normalized head coefficients $\alpha_{h}=\text{Softmax}(\mathbf{v})_{h}$, yielding the final fused distribution:
$\hat{P}=\sum_{h=1}^{H}\alpha_{h}\cdot O_{h},\quad\hat{P}\in\Delta^{C}$ (3)
This formulation ensures probability coherence while supporting flexible integration of heterogeneous predictors. It further enables interpretability through learned weighting patterns across both modalities and classes.
To combine modalities at the feature level, a simple lightweight fusion model is proposed. Each modality is projected into an eight-dimensional emotion space via a FCL. These logits are concatenated and processed by a shallow network with FCLs that captures inter-modality dependencies and outputs final emotion predictions.
@@ 4 Experiments
14 Scene+Label text & CLIP & Cosine similarity 10.84 13.09 11.97 11.79 30.41 24.10
@@
15 Models ID 1 & 5 & 6 & 10 & 14 & PF 44.92 49.58 47.25 47.01 46.89 46.95
16 Models ID 1 & 5 & 6 & 10 & 13 & MHPF 45.21 50.54 47.88 46.41 46.47 46.44
17 Models ID 1 & 5 & 6 & 10 & 13 & PF 46.95 51.08 49.02 45.47 46.00 45.74
@@
Table 2: Performance comparison (F1, %) of the SOTA methods.
Method Year Modality C-EXPR-DB
Wang et al. [32] 2024 V 18.45
Ryumina et al. [27] 2024 AV 22.01
Yu et al. [34] 2024 V 22.40
Liu et al. [19] 2025 V 22.81
Ryumina et al. [26] 2024 AV 25.56
Richet et al. [24] 2025 AVT 25.91
Savchenko [28] 2024 AV 27.08
Savchenko [29] 2025 V 32.43
Qiu et al. [22] 2024 AV 55.26
Lu et al. [20] 2025 V 57.23
Liu et al. [18] 2025 V 60.63
Ours 2025 V 34.86
@@ 4.2 Experimental Results
The experimental results demonstrate (in Table 1) varying performance across different model configurations on three benchmark datasets: Aff-Wild2 (8 classes), AFEW (8 classes), and C-EXPR-DB (7 classes). Individual models utilizing visual features show moderate effectiveness, with combinations of face-based features and EmoAffectNet achieving 34.08 average performance on Aff-Wild2 and 45.80 on AFEW. When enhanced with temporal modeling using either Transformer or Mamba architectures, these models exhibit divergent behavior: the Transformer variant improves Aff-Wild2 performance to 40.74 but degrades AFEW results to 37.01, while Mamba achieves more balanced gains of 38.29 and 37.98 respectively. Similarly, CLIP-based visual representations combined with Mamba yield competitive scores of 42.98 (Aff-Wild2) and 37.46 (AFEW), outperforming their Transformer counterparts.
Models incorporating scene-level features and language encoders present mixed outcomes. Among them, the combination of scene features with RoBERTa embeddings achieves 40.49 on Aff-Wild2 and 32.17 on AFEW, along with notable performance on C-EXPR-DB [15] (macro-F1 of 34.62 by pair wise sum method and macro-F1 of 34.85 by feature similarity method), suggesting its potential for capturing contextual cues. In contrast, Jina-embeddings perform slightly weaker overall but maintain stable recognition rates across both emotion corpora. Audio-based models generally underperform compared to visual and textual modalities, with WavLM and U-Net-aware frontends yielding average scores around 33.67 and 29.58 on Aff-Wild2, respectively. Text-based models, particularly those leveraging CLIP embeddings, achieve relatively high Aff-Wild2 scores (up to 40.83), although they significantly drop on AFEW, indicating domain mismatch or limited linguistic expressiveness in that dataset.
Among multi-modal fusion strategies, the best models combine multiple feature modalities (visual, textual, and audio) alongside advanced fusion techniques such as simply Probability Fusion (PF) (MHPF with one head) and MHPF (with three heads). The top configuration employing PF on models 1, 5, 6, 10, 14, and 15 achieves the highest score on Aff-Wild2 (49.02 average), while another variant of MHPF on the same set of models yields the best performance on AFEW (47.53 average). Notably, this ensemble also attains a macro-F1 of 27.08 on C-EXPR-DB [16], demonstrating its generalization capability across domains and annotation schemes. These results highlight the importance of complementary signal integration and adaptive fusion mechanisms in achieving robust CE recognition performance under challenging real-world conditions.
Despite the gain on emotional corpora due to multimodal fusion, on the C-EXPR-DB [9] corpus CER, the best performance is achieved by the scene-based model. This result shows that with correct prompting for vision-language model, reliable results can be achieved for both basic emotion recognition and CER. Comparison with the SOTA results (Table 2) shows that the proposed approach outperforms all the approaches that were not trained on the target task, losing only to the approaches [22, 20, 18], which do not represent a zero-shot solution.

### v2

@@ 1 Introduction
Our approach introduces a novel multimodal framework specifically designed for CER. It integrates six diverse modalities – static and dynamic facial expressions, audio signals, scene context, textual descriptions, scene-label alignment, probability and feature fusion – into a unified system. In contrast to prior work that relies primarily on static facial features [20], our method incorporates temporal dynamics using Transformer- and Mamba-based architectures, enabling more accurate modeling of affective evolution over time.
Moreover, unlike systems that require extensive fine-tuning on labeled CE corpora [17, 16], our pipeline leverages zero-shot-capable components. These include Contrastive Language-Image Pretraining (CLIP)-driven label matching and semantic scene analysis using Qwen-VL, allowing the model to generalize beyond predefined emotion classes without domain-specific retraining. We also propose a Multi-Head Probability Fusion (MHPF) model that adaptively balances contributions from individual modality-specific predictors.
@@ 2 Related Work
We compare our proposed method with the State-of-the-Art (SOTA) methods presented in the scope of the 7th, 8th Affective Behavior Analysis in-the-Wild (ABAW) Competitions [12]. This challenge has seen the development of several CER systems, including our own [22].
Savchenko [23] proposed a visual-only method using lightweight EmotiNet models pre-trained for basic facial expression recognition. CE predictions were obtained by aggregating the outputs of basic emotions and applying temporal smoothing with box and Gaussian filters. This method enhanced stability while maintaining computational efficiency, making it applicable to real-time scenarios.
@@
In the methods proposed by [17, 16], additional corpora annotated with CEs were used to train models. Therefore, our method is more comparable to those of [20, 23], which rely on pretrained components and avoid direct training on the target corpus, thus operating in a zero-shot or lightly adapted setting.
@@ 3.1 Face Models
For each segment, the frame rate is downsampled to 20 frames per 4-second window, ensuring uniform temporal resolution and reducing computational complexity. Facial detection is performed using a two-stage method: an initial YOLO v11222https://github.com/akanametov/yolo-face model, pre-trained on WIDER FACE, and fine-tuned for real-world conditions, detects faces under challenging conditions, while a secondary refinement step based on MediaPipe FaceMesh [18] validates face geometry and filters out false positives detections.
@@ 4 Experiments
14 Scene-Label text & CLIP & Cosine similarity 10.84 13.09 11.97 11.79 30.41 24.10
@@
15 Models ID 1 & 5 & 6 & 10 & 14 & MHPF 45.21 50.54 47.88 46.41 46.47 46.44
@@
Table 2: Performance comparison (F1, %) of the SOTA methods. V is video modality, AVT is audio, video, and text modalities. ZH is zero-shot setup. FT is fine-tuned models on target task.
Method Setup Modality C-EXPR-DB
Liu et al. [17] FT V 22.81
Richet et al. [20] ZH AVT 25.91
Savchenko [23] ZH V 32.43
Liu et al. [16] FT V 60.63
Ours ZH V 34.86
@@ 4.2 Experimental Results
The experimental results demonstrate (in Table 1) varying performance across different model configurations on three benchmark datasets: AffWild2 (8 classes), AFEW (8 classes), and C-EXPR-DB (7 classes). Individual models utilizing visual features show moderate effectiveness, with combinations of face-based features and EmoAffectNet achieving 34.08 average performance on AffWild2 and 45.80 on AFEW. When enhanced with temporal modeling using either Transformer or Mamba architectures, these models exhibit divergent behavior: the Transformer variant improves AffWild2 performance to 40.74 but degrades AFEW results to 37.01, while Mamba achieves more balanced gains of 38.29 and 37.98 respectively. Similarly, CLIP-based visual representations combined with Mamba yield competitive scores of 42.98 (AffWild2) and 37.46 (AFEW), outperforming their Transformer counterparts.
Models incorporating scene-level features and language encoders present mixed outcomes. Among them, the combination of scene features with RoBERTa embeddings achieves 40.49 on AffWild2 and 32.17 on AFEW, along with notable performance on C-EXPR-DB [13] (macro-F1 of 34.62 by pair wise sum method and macro-F1 of 34.85 by feature similarity method), suggesting its potential for capturing contextual cues. In contrast, Jina-embeddings perform slightly weaker overall but maintain stable recognition rates across both emotion corpora. Audio-based models generally underperform compared to visual and textual modalities, with WavLM and U-Net-aware frontends yielding average scores around 33.67 and 29.58 on AffWild2, respectively. Text-based models, particularly those leveraging CLIP embeddings, achieve relatively high AffWild2 scores (up to 40.83), although they significantly drop on AFEW, indicating domain mismatch or limited linguistic expressiveness in that dataset.
The best performing multimodal configuration employing MHPF on models ID 1, 5, 6, 10, 13, and 14 achieves the highest scores on AffWild2 (48.77 average two metrics) and AFEW (47.34 average). Notably, this ensemble also attains a macro-F1 of 27.08 on C-EXPR-DB [14], demonstrating its generalization capability across domains and annotation schemes. Despite the gain on emotional corpora due to multimodal fusion, on the C-EXPR-DB [15] corpus CER, the best performance (macro-F1 = 34.85) is achieved by the scene-based model. This result shows that with correct prompting for a vision-language model, reliable results can be achieved for both basic emotion recognition and CER. Comparison with the SOTA results (Table 2) shows that the proposed approach outperforms all the approaches [17, 23] that were not trained on the target task, losing only to the approaches [16], which do not represent a zero-shot solution.